What is the quadratic form for g(x)=2(4x+1)²?
Answer by sdmmadam@yahoo.com(530)

What is the quadratic form for g(x)=2(4x+1)²?

g(x)=2(4x+1)²
= 2[(4x)^2 + 2X(4x)X (1) + (1^2)]
(using (a+b)^2 = (a^2) +2(a)(b) + (b^2) and here a = 4x and b = 1)
=2X(16x^2+8x+1)
=32x^2+16x+2
which is the desired quadratic form.